Circular No. 8296

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004ac AND 2004ad M. Armstrong, Rolvenden, U.K., reports the discovery of an apparent

supernova (mag 17.0) on unfiltered CCD images taken on Mar. 1.796 and 2.789 UT. The new object is located at α = 2^{h}02^{m}26^{s}.39, δ = +32^{o}04′22′′.2 (equinox 2000.0), which is ≈ 4′′ east and 1′′.6 north of the center of NGC 789. Nothing was visible at this position on Armstrong’s images taken on 2003 Sept. 24 (limiting mag 19.5), 2004 Jan. 9 (limiting mag 18.5), and Feb. 9 (limiting mag 19.0) or on Palomar Sky survey red and blue plates.

Corrigendum. On IAUC 8295, ‘Supernova 2004ac in IC 4769’, line 5, for 9′′ west read 2′′ west

SUPERNOVAE 2004ae, 2004af, 2004ag, 2004ah, 2004ai, 2004aj M. Riello, S. Benetti, E. Cappellaro, F. Patat, M. T. Botticella, G.

Altavilla, A. Pastorello, M. Turatto, and S. Valenti report the discovery of six supernovae on CCD frames taken on Feb. 18 during a search performed with the Max-Planck-Institut 2.2-m telescope (+ Wide Field Imager) at the European Southern Observatory (ESO).

SN  α_{2000}  δ_{2000}  V  Offset  z Type
2004ae  4^{h}28^{m}17.^{s}89 −36^{o}18′55′′.0 23.3 < 0′′.3  0.480 II
2004af  5 38 03.91 −23 59 00.2 22.5 3′′.2 W, 4′′.3 S 0.056 Ic
2004ag 9 51 01.49 −20 50 37.5 23.0 < 0′′.5 0.362 II
2004ah 10 45 47.41 − 0 06 58.1 23.3 < 0′′.8  0.480 Ia
2004ai 13 54 26.09 −12 41 15.9 23.4 1′′.0 S  0.590 Ic
2004aj 14 20 37.58 −12 24 14.4 22.3  0.247 Ia

Nothing was visible at the above positions on reference images taken during 2001–2002. Spectroscopy (range 600–1100 nm), obtained on Feb. 26 and 27 with the ESO 8.2-m YEPUN telescope (+ FORS2), reveals that each supernova was caught around maximum, with the exception of 2004af, which shows very strong nebular emission lines of Mg I], Na I, [O I], [Ca II], and Ca II, and 2004af bears a strong resemblance with that of SN 1998bw at ∼ 6 months past maximum light (Patat et al. 2001, Ap.J. 555, 900). The enhanced emission that is observed in the range 400–550 nm may also be attributed to a blend of Fe lines. The absolute magnitude of 2004af, M(R) = −14.3 (H_{o} = 72 km/s/Mpc), is unusually bright for a type-Ib/c event at these phases and implies a maximum brightness M(R) < −19. This, coupled with the spectral appearance, make 2004af a hypernova candidate.

2004 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green